OpenStudy (anonymous):

evaluate integral(D) xcosy dA using fubini's theorem, where D is the region bounded by y=0, y=x^2 and x=1

OpenStudy (anonymous):

What's is Fubini's theorem?

OpenStudy (anonymous):

if D is the type 1 region such that [D= {(x,y)/a<=x<=b, g1(x)<=y<=g2(x)} with g1(x) and g2(x) continuously differentiabl functions on [a,b] and if f:D\[\rightarrow\]R is continuous then \[\int\limits_{}^{}\int\limits_{}^{} f(x.y)dA = \int\limits_{a}^{b}\int\limits_{g1(x)}^{g2(x)} f(x,y) dydx\]